# Jiaying Wu

### v1

@@ Abstract
Synthetic face recognition (SFR) aims to generate synthetic face datasets that mimic the distribution of real face data, which allows for training face recognition models in a privacy-preserving manner. Despite the remarkable potential of diffusion models in image generation, current diffusion-based SFR models struggle with generalization to real-world faces. To address this limitation, we outline three key objectives for SFR: (1) promoting diversity across identities (inter-class diversity), (2) ensuring diversity within each identity by injecting various facial attributes (intra-class diversity), and (3) maintaining identity consistency within each identity group (intra-class identity preservation). Inspired by these goals, we introduce a diffusion-fueled SFR model termed $\text{ID}^{3}$. $\text{ID}^{3}$ employs an ID-preserving loss to generate diverse yet identity-consistent facial appearances. Theoretically, we show that minimizing this loss is equivalent to maximizing the lower bound of an adjusted conditional log-likelihood over ID-preserving data. This equivalence motivates an ID-preserving sampling algorithm, which operates over an adjusted gradient vector field, enabling the generation of fake face recognition datasets that approximate the distribution of real-world faces. Extensive experiments across five challenging benchmarks validate the advantages of $\text{ID}^{3}$.
**footnotetext: Equal first authors. The order was determined by numpy.random.rand().
@@ 4.2 Implementation Details
IDiff-Face FFHQ 97.10 82.00 76.65 78.40 86.32 84.09
@@ Appendix A Appendix A
Note that these reasonable assumptions are applied in (Xu et al., 2023; Li et al., 2021; Hasnat et al., 2017). Now we can specify the value of the scalar $C$:

### v2

@@ Abstract
Synthetic face recognition (SFR) aims to generate synthetic face datasets that mimic the distribution of real face data, which allows for training face recognition models in a privacy-preserving manner. Despite the remarkable potential of diffusion models in image generation, current diffusion-based SFR models struggle with generalization to real-world faces. To address this limitation, we outline three key objectives for SFR: (1) promoting diversity across identities (inter-class diversity), (2) ensuring diversity within each identity by injecting various facial attributes (intra-class diversity), and (3) maintaining identity consistency within each identity group (intra-class identity preservation). Inspired by these goals, we introduce a diffusion-fueled SFR model termed $\text{ID}^{3}$. $\text{ID}^{3}$ employs an ID-preserving loss to generate diverse yet identity-consistent facial appearances. Theoretically, we show that minimizing this loss is equivalent to maximizing the lower bound of an adjusted conditional log-likelihood over ID-preserving data. This equivalence motivates an ID-preserving sampling algorithm, which operates over an adjusted gradient vector field, enabling the generation of fake face recognition datasets that approximate the distribution of real-world faces. Extensive experiments across five challenging benchmarks validate the advantages of $\text{ID}^{3}$. Code is released at: https://github.com/hitspring2015/ID3-SFR.
@@ 4.2 Implementation Details
IDiff-Face***Equal first authors. The order was determined by numpy.random.rand(). FFHQ 97.10 82.00 76.65 78.40 86.32 84.09
@@
**footnotetext: Replicated by using its official implementation: https://github.com/fdbtrs/IDiff-Face.
@@ Appendix A Appendix A
Note that these reasonable assumptions are also held in (Xu et al., 2023; Li et al., 2021; Hasnat et al., 2017). Now we can specify the value of the scalar $C$:
@@ Appendix D Related Work
Latent Diffusion Models. There exist many diffusion-based models (LDMs) (e.g., Face0 (Valevski et al., 2023), PhotoMaker (Li et al., ), FaceStudio (Yan et al., 2023), InstantID (Wang et al., 2024)) which use ID attributes to assist in generating images. However, we note that these LDMs are designed for image generation but not for SFR. Our empirical findings further suggests these LDMs do not perform reasonably well even when applied to SFR. We found that although these LDMs claim to be ID-preserving in the pixel space, their feature embeddings are not discriminative enough for face recognition, since there is no inductive bias (neither loss functions nor architectures) to achieve face discriminativeness.
@@ Appendix E Ablation Study (ii)
Appendix F Attribute Analysis of Generated Datasets
In this section, we perform an attribute analysis of the generated face datasets by ID3. As shown in Figure A.3 and Figure A.4, from the distribution, we observe that the highest age in our training set is 70; the largest pose is 60 in degree and the smallest -60 in degree. Our model can interpolate within these ranges but is less likely to extrapolate outside of these ranges. To generate face images of large pose and high age, one can collect more such data and add them to the training dataset, which increases their occurrences during training.
To examine whether the attributes we used in our paper are fit for SFR tasks, we perform the following ablation study: as we increase intra-class pose variation, the SFR performances on cross-pose test sets (including CFPFP and CPLFW) are boosted whereas the performances on cross-age test sets (including AgeDB and CALFW) remain almost unchanged. Results and distribution plots are shown in Table A.1, Figure A.3 and Figure A.4. From these results, we observe that the distribution of pose angle on FFHQ, Dataset by IDiffFace, Dataset by ID PoseOnly 1 and 2 are all unimodal. And their performances are inferior to Dataset by ID PoseOnly 3 which exhibits multimodal distribution of pose angle.
Figure A.3: Pose distribution of datasets generated by different models.
Figure A.4: Pose distribution (pitch, yaw, roll) and age distribution on FFHQ
Table A.1: Datasets generated by different models (Column 1), the attribute statistics for each dataset (Column 2, 3), and the FR performance of FR models trained on them, respectively (Column 4-8).
Pose Mean Pose Var LFW CFP-FP CPLFW AgeDB CALFW
FFHQ 11.44 233.97
IDiffFace 11.62 222.62 97.10 82.00 76.65 78.40 86.32
ID3 PoseOnly 1 8.21 109.76 95.33 78.41 73.48 79.76 86.03
ID3 PoseOnly 2 9.02 110.18 95.58 80.91 73.60 79.45 85.93
ID3 PoseOnly 3 14.14 247.05 95.83 82.87 75.77 79.45 86.90